OpenStudy (anonymous):

Solve the following word problem by using an equation and factoring. The sum of a positive number and its square is 20. What is the number? A. 1 B. 2 C. 3 D. 4 E. 5

OpenStudy (mertsj):

x+x^2=20 x^2+x-20=0 (x+5)(x-4)=0 x=-5 or x = 4 Discard -5 because the problem calls for a positive number.